Hi I have this due tonight and I just do not understand can you help me out please? Thank You-
Write a system of two equations in two unknowns for each problem. Solve each system by substitution.
Household Income. Alkena and Hsu together earn $84,326 per year. If Alkena earns $12,468 more per year than Hsu, then how much does each of them earn per year?

Answer by gokhalesujata(4)

let alkenas salary be $x
let Hsus salary be $y
fact one x+y=84,326
also x=y+12,468 fact two
substituting x=y+12468 in equation x+y=84326 we get
y+12468+y=84326
2*y=84326-12468
2*y=71858
y=71858/2=35929
x=35929+12468=48397